Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A jar contains 6 red marbles and 9 blue marbles. If Evelyn

Expert replies
by swerve » Mon May 21, 2018 10:22 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A jar contains 6 red marbles and 9 blue marbles. If Evelyn reaches into the jar and simultaneously draws two marbles at random, what is the probability that she will draw two marbles of the same color?

A. 2/7
B. 12/35
C. 3/7
D. 17/35
E. 25/35

The OA is D.

Please, can anyone explain this PS question for me? I tried to solve it but I can't get the correct answer. I need your help. Thanks.
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon May 21, 2018 10:39 am
swerve wrote:A jar contains 6 red marbles and 9 blue marbles. If Evelyn reaches into the jar and simultaneously draws two marbles at random, what is the probability that she will draw two marbles of the same color?

A. 2/7
B. 12/35
C. 3/7
D. 17/35
E. 25/35
Case 1: 2 green marbles
P(1st marble is green) = 6/15. (Of the 15 marbles, 6 are green.)
P(2nd marble is green) = 5/14. (Of the 14 remaining marbles, 5 are green.)
To combine these probabilities, we multiply:
6/15 * 5/14 = 1/7.

Case 2: 2 red marbles
P(1st marble is red) = 9/15. (Of the 15 marbles, 9 are red.)
P(2nd marble is red) = 8/14. (Of the 14 remaining marbles, 8 are red.)
To combine these probabilities, we multiply:
9/15 * 8/14 = 12/35.

Since a good outcome will be yielded by Case 1 or Case 2, we ADD the fractions in blue:
1/7 + 12/35 = 5/35 + 12/35 = 17/35.

The correct answer is D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

swerve wrote:A jar contains 6 red marbles and 9 blue marbles. If Evelyn reaches into the jar and simultaneously draws two marbles at random, what is the probability that she will draw two marbles of the same color?

A. 2/7
B. 12/35
C. 3/7
D. 17/35
E. 25/35
P(both are same color) = P(1st marble is red AND 2nd marble is red OR 1st marble is blue AND 2nd marble is blue)
= [P(1st marble is red) x P(2nd marble is red)] + [P(1st marble is blue) x P(2nd marble is blue)]
= [6/15 x 5/14] + [9/15 x 8/14]
= [2/5 x 5/14] + [3/5 x 8/14]
= 10/70 + 24/70
= 34/70
= 17/35

Answer: D

Cheers,
Brent
Last edited by Brent@GMATPrepNow on Mon May 21, 2018 10:44 am, edited 1 time in total.
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Sionainn@PrincetonReview » Mon May 21, 2018 10:43 am
There are two outcomes that meet the requirement either BB OR RR. When working with probability it boils down to putting want you want over the total number. When you have multiple fractions, typically AND means you need to multiply the fractions and OR means you need to add. So for this problem we can translate it into the following:

Red AND Red OR Blue AND Blue

Now plug values into this, keeping in mind on the second draw of each option you will have fewer marbles available, so this gives you

6/16*5/14 + 9/15*8/14

Cancel out common factors to make the numbers easier to work with.

1/1*1/7+3/5*4/7

Multiply and then add.

1/7 + 12/35 = 17/35.

So the answer is choice D.

Take care,
Sionainn
BA - Stanford University, MPP - Harvard University
Instructor, tutor for Princeton Review and Airbnb host
In other words a blend of Jamie Escalante from Stand and Deliver, Julie from The Love Boat, and Schneider the Super from One Day at a Time.
Image
Curious How You’ll Score? Take a FREE GMAT® practice test or sample class
Ready to Prep? Exclusive discounts for Beat The GMAT® members HERE
Join the discussion

by regor60 » Tue May 22, 2018 6:45 am
The total number of ways to withdraw 2 marbles regardless of color is 15!/2!13!

The number of ways to draw two red marbles is 6!/2!4!

The number of ways to draw two blue marbles is 9!/2!7!

The probability of drawing either one of the above is (6!/2!4! + 9!/2!7!)/(15!/2!13!) = [spoiler]17/35, D[/spoiler]
Join the discussion